OpenStudy (anonymous):

C = 5/9 (F-32) Solve for F show all steps

OpenStudy (anonymous):

C = 5/9 (F-32) C = (F-32) 5/9 * c= f-32 5/9 * c + 32= f

OpenStudy (anonymous):

oops 9/5 * c +32 =f

OpenStudy (anonymous):

C = 5/9 (F-32) 9/5 *c = f-32 multiply both sides by 9/5 to get rid of 5/9 on rhs 9/5 *c + 32 = f add 32 to both sides to get rid of -32 on rhs

OpenStudy (anonymous):

so what do i put

OpenStudy (anonymous):

f= 9/5 *c + 32

OpenStudy (anonymous):

Timo she used you! lol

OpenStudy (anonymous):

this is baby math I barely got used :p

OpenStudy (anonymous):

lol